a, b, n ∈ Z with n > 1. Prove that if a ≡ b (mod n) then 2a ≡ 2b (mod n).
Answer by rothauserc(4718)

a congruent to b modulus n means that a-b is divisible by n
:
then we have
:
a - b = nc where c is an integer
:
then we have
:
a = b + nc
:
multiply both sides of the = by 2
:
2a = 2b + 2nc
:
we have a, b, n, c are integers, then
:
2c is an integer say p, then
:
***********************
2a = 2b + np = 2b mod n
We are done
***********************
: